# The Sloan Digital Sky Survey Reverberation Mapping Project: Insights on Maximizing Efficiency in Lag Measurements and Black-Hole Masses

### v1

@@ 1 Introduction
Recently, industrial-scale RM programs such as the Sloan Digital Sky Survey Reverberation Mapping (SDSS-RM) project (Shen et al., 2015a) and the Australian Dark Energy Survey RM (OzDES) project (King et al., 2015) have started to use multi-object spectrographs on survey telescopes to carry out RM observations of large samples of diverse quasars. While these surveys have been successful, a significant fraction of the sample lacks well-detected lags. This is because extracting reliable and well-defined lag measurements from industrial-scale RM surveys is challenging (see § 3 for more details). This is further complicated because, at higher redshifts, observations are limited to more luminous quasars, which tend to have longer time delays due to the combined effects of cosmological time dilation and the canonical relation (Bentz et al., 2013). Additionally, the inherent lower variability amplitude observed in more luminous quasars (e.g.,Vanden Berk et al., 2004; MacLeod et al., 2010) further reduces the probability of successful lag recovery. Industrial-scale RM faces challenges beyond observational costs due to the massive volume of light curves. Efficiently analyzing these light curves and selecting the most significant and reliable lag measurements is computationally intensive. As future surveys like the Legacy Survey of Space and Time (LSST; e.g., Ivezić et al. 2019) produce millions of light curves, developing strategic analysis tools will be crucial for timely and effective RM studies. Although this study focuses on emission-line RM, the methodology developed here can be readily extended to investigate continuum RM, provided that a sufficient number of continuum RM lag measurements are available.
@@
In large-scale surveys like the SDSS-RM, a comprehensive understanding of the interplay between the basic quasar properties, the quasar intrinsic variability characteristics, the survey design, and the observational sensitivity is critical for evaluating and forecasting the program’s success in terms of lag-measurement yield and limitations. The detection of reverberation lags relies heavily upon the monitoring program’s design, specifically the observing cadence, total observing baseline, presence and distribution of seasonal/weather gaps, and the signal-to-noise ratio (S/N) of the flux measurements. Previous light-curve simulations of the SDSS-RM survey have already studied the correlation between time-series analysis methods and survey yields (Li et al., 2019). The OzDES campaign has similarly employed light-curve simulations replicating source variability, measurement errors, and observing cadence to assess lag measurement reliability criteria (Penton et al., 2022). However, there has never been a comprehensive assessment connecting physical sample properties and successful lag measurement. The present work focuses on a sub-sample of the most reliable RM lag measurements from the SDSS-RM survey, specifically those quasars exhibiting well-detected lag measurements (i.e., the “gold sample”, see § 3 for more details). By investigating the physical and statistical properties of this sub-sample, we aim to identify the characteristics of the already observed quasars that are most favorable for efficient RM lag measurements from the SDSS-RM survey. We also explore the implications of a reduced cadence in the SDSS-RM survey, enabling more efficient target selection, cadence design, and time-series analysis for future RM campaigns like the 4MOST survey’s TiDES program (de Jong et al., 2019; Swann et al., 2019), which aims to conduct RM observations of 700 quasars over five years, continuing to monitor some of the previously studied RM fields.
@@ 4.1 Continuum Luminosity
Intrinsic AGN luminosity is considered the primary driver of reverberation lag times. More luminous quasars tend to have larger time lags as observed by the relation (Bentz et al., 2013; Fonseca Alvarez et al., 2019), which matches basic photoionization expectations (). Furthermore, luminous quasars often exhibit lower amplitude variability (MacLeod et al., 2012; Vanden Berk et al., 2004), potentially reducing the likelihood of detecting a reliable lag measurement within short observing campaigns. To explore this connection, we adopt the luminosity as a measure of the quasar continuum luminosity. Whenever possible, we utilize directly measured values. For sources lacking direct measurements, we reconstruct them using , , and the bolometric corrections from Richards et al. (2006). Figure 1 illustrates the redshift and luminosity distribution of our targets. We performed a global comparison of the continuum luminosity within the gold samples of early SDSS-RM studies and the current 7-year lag measurements (Grier et al., 2017, 2019; Homayouni et al., 2020; Shen et al., 2024). This analysis did not reveal any statistically significant changes in the continuum luminosity distribution over the multi-year analysis. Additionally, we find no dependence between lag measurement success and source luminosity. The median continuum luminosity of the gold sample (in logarithmic scale) is , , and for H, Mg II, and C IV respectively for the early SDSS-RM lag measurements, which is similar to the median range of , , and in the parent sample of each targeted emission line. Here the uncertaintly on the median vales are computed through bootstrap method. We find similar ranges for the gold-lag measurements in the 7-year results of Shen et al. (2024).
@@ 4.2 Rest-frame Equivalent Width
The rest-frame equivalent width (REW) of emission lines is another potential factor influencing successful lag measurement success. We investigate whether targets with larger REWs correlate with the most reliable lag measurements. For each emission line, we examine the potential relationship between its REW using PrepSpec outputs. REWs are measured from the continuum and flux provided by PrepSpec, utilizing either one-year, four-year, or seven-year average spectra. Figure 2 shows the comparisons between early-year and 7-year SDSS-RM data for the same emission lines, which reveal relatively consistent REWs over time. Furthermore, we find no direct correlation between the fraction of quasars with gold-lag measurements. Figure 2 depicts how the gold samples in each study are uniformly distributed within their respective parent samples, and the gold-lag success fraction does not reveal any specific trends with REW. While the one-year and four-year PrepSpec spectral fits did not incorporate any Fe II emission lines, the seven-year data included an Fe II template. Consequently, the 7-year measurements of Shen et al. 2024 (blue data points in Figure 2) appear to be systematically offset above the 1:1 line, suggesting that for all the three emission lines, the gold-lag estimates may be biased towards larger REWs. However, we did not observe a higher success rate for the gold-lag sample, except for C IV.
@@ 5.1 Fractional Variability
The results of this investigation are presented in Figure 4. We further compared the fractional rms variability between the early-year and 7-year studies for each emission line bottom panels in Figure 4. It is noteworthy that the 4-year and 7-year results largely follow a 1:1 trend. The deviation observed in the comparison between the 1-year and 7-year H fractional rms variability (bottom left corner of Figure 4) can be attributed to the variability of quasars on the observed-frame timescales. The damping timescale of quasars are longer than the seasonal monitoring duration.
@@ 5.3 Durbin-Watson Statistic
While our findings suggest the importance of variability as measured by SNR2, instances exist where a gold lag remains undetected despite relatively high SNR2 values. For time-series analysis software to effectively detect an RM lag, the light curve often requires a distinct “hook” or inflection feature in the continuum and line light curves. In this section, we investigate whether the presence of serial correlations in the continuum and emission line light curve would lead to a higher success in detecting a gold-lag measurement. Serial correlation refers to the dependence of current residuals (from a regression model) on past residuals, indicating a patterned behavior in the time series. A commonly used test for first-order autocorrelation, which assumes independent error terms, is the Durbin-Watson test (Durbin & Watson 1950). The Durbin–Watson test statistic is given by
@@ 6.1 Cadence Intensity in SDSS-RM
Finally, we run PyROA for the simulated line light curves with their original continuum light curves and apply lag identification and alias removal process to get the final lag measurement. The reason we keep the original continuum light curves and do not remove epochs from them is that the photometric observations for continuum light curves are not as expensive to make as the spectroscopic observations and are made by telescopes worldwide, so the continuum light curves are therefore densely sampled. Specifically, we have 800 epochs along 11 years and 158 epochs in 2014 (the first year of line light curves) on average, which is extremely dense with a cadence of 3 days per epoch approximately. We also perform some tests on running PyROA for cadence reduced light curves for both continuum and emission lines. The distribution of interval between one epoch to the next is almost uniform for each year, i.e., we just reduced 20% of the total epochs for continuum light curves of several targets selected randomly from the gold sample and found similar lag PDFs using PYROA Therefore, for convenience, removing epochs in the first year in the following texts refers to randomly reducing the density of epochs in the first year of the original emission line light curve to 40% unless otherwise specified.
@@ 6.2 Impact of Reduced Cadence on Lag Success
We run all the simulated light curves through the same lag identification pipeline as Shen et al. (2024) and perform the following statistical analyses. We use three criteria to assess the difference between the original lag measurements and the simulated lag measurements for targets from the gold sample, and display the variation of simulated lags in Figure 8. The three criteria are:
@@
We determine the relative difference between the median of the 50 simulated lag and the original lag as , and require the relative difference less than 34%.
The distributions of the lag selection criteria, , and are displayed on the left panels in Figure 8. Most of the lag posterior samples given by PyROA will still have a single, prominent peak in the lower-cadence simulations, so we do not show the changes for the lag selection criteria . While the statistical values for and generally decreased in the lower-cadence simulations, most simulated lags still met the significance criteria. Figure 8 highlights several targets with original and values near the selection threshold that resulted in insignificant lag measurements under reduced cadence conditions. The simulated lags that were initially considered insignificant but later became significant are primarily found in cases with lower overall light curve quality. This suggests that the reduction in measurement quality during the first season pushed these lags below the original selection criteria. Despite having the lowest median observed-frame lag (31 days), the H sample still has the highest ratio of significant results. This may be attributed to the generally higher lag and values in the H sample, indicating its robustness to epoch reduction.
@@
We display the consistency level for significant results in the middle column of Figure 8. As described above, the can quantitatively describe how well the original lag and the simulated lag are consistent within a given margin of error. We divided the sample into long- and short-lag groups based on their median original lags: 31 days for H, 155 days for Mg II, and 140 days for C IV. Short-lag targets exhibited a higher degree of consistency () compared to long-lag targets across all emission lines. The most pronounced difference was seen for H, with short-lag targets achieving a 92% 1% consistency rate, surpassing the 79% 1% of long-lag targets. Mg II and C IV also exhibited similar trends, with short-lag targets outperforming long-lag targets by small margins. However, among the three emission lines, C IV (with the longest median original lag) demonstrated the highest overall ratio of simulations that maintained consistent results within 1 errors, while H and Mg II exhibited lower ratios. The short-lag sample may exhibit higher consistency due to its smaller absolute changes in lags and associated errors compared to the long-lag sample.
Figure 8 (right column) illustrates the distribution of relative differences between simulated and original lags for significant results of H, Mg II, and C IV. The relative difference, , quantifies the change in the median simulated lag relative to the original. Long-lag targets consistently exhibited a higher proportion of “similar” results () compared to short-lag targets. For H, this ratio was 90% 1% for long lags and 85% 1% for short lags. Similar trends were observed for Mg II (94% vs. 86%) and C IV (95% vs. 89%). However, if we require both statistical significance and consistency within , the lag recovery ratios are 81% for H, 76% for Mg II, and 86% for C IV. Targets with longer original lags are slightly more susceptible to a 40% reduction in cadence during the first year (13 epochs).
To inform future cadence design, we examined the relationship between lag recovery ratio, -band magnitude, and redshift, given that SDSS-RM targets are selected based on flux limits. Figure 9 presents an overview of the lag recovery ratio binned by redshift (bin size = 0.4) and -band magnitude (bin size = 1). Overall, we observed a trend of higher success ratios for brighter -band magnitudes and lower redshifts. Several bins, particularly those near the boundary of the redshift-magnitude parameter space, contain fewer than five targets and may not provide statistically meaningful results. We therefore excluded these bins from Figure 9. Given our results based on seven years of observations from SDSS-RM, the limited number of significant lags detected at these redshifts and magnitudes, as also indicated by mock data in Shen et al. (2015a), underscores the challenges of measuring significant lags at these cosmological distances and magnitudes.
@@ 7 Summary and Implications for Future Industrial-Scale Surveys
The Durbin-Watson () statistic is a reliable indicator of gold-lag measurements, particularly when applied to emission line light curves. While the statistic for the continuum light curve shows a weaker correlation, analyzing the emission line reveals a stronger trend. For Mg II, the statistic exhibits a weaker correlation with gold lag success compared to H and C IV. This is likely due to Mg II’s lower variability amplitude (Sun et al., 2015). Overall, our analysis suggests that the statistic, especially when applied to emission line light curves, is a valuable tool for efficiently filtering targets and predicting the likelihood of obtaining reliable lag measurements, particularly in large datasets.
Additionally, we investigate the impact of reducing the SDSS-RM cadence on lag measurements based on the observed quasar light curves of Shen et al. (2024). We simulate multi-year light curves with 40% reduction in cadence during the first year and compare with the existing lag measurements for the same target. While most statistically significant lags are preserved after cadence removal (90% or higher for all emission lines studied: H, Mg II, and C IV), a small percentage become insignificant. The consistency between original and simulated lags (within a margin of error) is generally higher for shorter lags compared to longer lags. However, when we consider the lag similarity, it is worse for shorter lags. This is likely due to the decreased temporal resolution, which can limit the ability to capture the rapid variations associated with shorter lags. Furthermore, we find that the recovery rate is slightly lower for fainter or higher redshift quasars (Figure 9). These results are consistent with earlier studies using shorter mock light curves Shen et al. (2015a). Overall, our simulations indicate that a mixed cadence strategy, featuring a period of higher cadence followed by lower cadence as in SDSS-RM, provides only a marginal improvement in gold-lag measurement compared to a more uniform cadence, which can still recover 90% of the lags.

### v2

@@ 1 Introduction
Recently, industrial-scale RM programs such as the Sloan Digital Sky Survey Reverberation Mapping (SDSS-RM) project (Shen et al., 2015a) and the Australian Dark Energy Survey RM (OzDES) project (King et al., 2015) have started to use multi-object spectrographs on survey telescopes to carry out RM observations of hundreds of diverse quasars. While these surveys have been successful, a significant fraction of the sample lacks well-detected lags. This is because extracting reliable and well-defined lag measurements from industrial-scale RM surveys is challenging (see § 3 for more details). This is further complicated because, at higher redshifts, observations are limited to more luminous quasars, which tend to have longer time delays due to the combined effects of cosmological time dilation and the canonical relation (Bentz et al., 2013). Additionally, the inherent lower variability amplitude observed in more luminous quasars (e.g.,Vanden Berk et al., 2004; MacLeod et al., 2010) further reduces the probability of successful lag recovery. Multi-object RM surveys typically target several hundred quasars, such as the 849 in SDSS-RM, 771 in OzDES, and the planned 700 in the 4MOST TiDES program. Industrial-scale RM faces challenges beyond observational costs due to the massive volume of light curves. These challenges primarily stem from the massive volume of data and the complexity of the analysis. Handling hundreds of quasar light curves, each with multiple emission lines, requires specialized software and expertise. Current methods for time-series analysis, outlier rejection, and lag-measurement validation, often involving manual inspection and rigorous statistical tests, are both time-consuming and computationally expensive. While the era of big data has advanced, the field of industrial-scale RM is still relatively young, and the optimization of lag-measurement pipelines is an ongoing challenge. Even with automated methods, human intervention remains essential for quality control and interpretation of results. Visual inspection of hundreds of light curves is a laborious task, requiring significant expertise and time. A typical light curve undergoes multiple stages of processing, including initial vetting, time-series analysis, significance testing, and potential visual inspection or statistical checks. This multi-step approach can extend the analysis timeline for large samples to several months. As future surveys like the Legacy Survey of Space and Time (LSST; e.g., Ivezić et al. 2019) produce millions of light curves, developing strategic analysis tools will be crucial for timely and effective RM studies. Although this study focuses on emission-line RM, the methodology developed here can be readily extended to investigate continuum RM, provided that a sufficient number of continuum RM lag measurements are available.
@@
In large-scale surveys like the SDSS-RM, a comprehensive understanding of the interplay between the basic quasar properties, the quasar intrinsic variability characteristics, the survey design, and the observational sensitivity is critical for evaluating and forecasting the program’s success in terms of lag-measurement yield and limitations. The detection of reverberation lags relies heavily upon the monitoring program’s design, specifically the observing cadence, total observing baseline, presence and distribution of seasonal/weather gaps, and the signal-to-noise ratio (S/N) of the flux measurements. Previous light-curve simulations of the SDSS-RM survey have already studied the correlation between time-series analysis methods and survey yields (Li et al., 2019). The OzDES campaign has similarly employed light-curve simulations replicating source variability, measurement errors, and observing cadence to assess lag measurement reliability criteria (Penton et al., 2022). However, there has never been a comprehensive assessment connecting physical sample properties and successful lag measurement. The present work focuses on a sub-sample of the most reliable RM lag measurements from the SDSS-RM survey, specifically those quasars exhibiting well-detected lag measurements (i.e., the “gold sample”, see § 3 for more details). By investigating the physical and statistical properties of this sub-sample, we aim to identify the characteristics of the already observed quasars that are most favorable for efficient RM lag measurements from the SDSS-RM survey. We also investigate how reducing the SDSS-RM cadence affects the number of gold-lag measurements. This analysis will inform future RM surveys, such as the 4MOST TiDES program (de Jong et al., 2019; Swann et al., 2019), by demonstrating the potential for achieving comparable lag results with less-frequent observations.
@@ 4.1 Continuum Luminosity
Intrinsic AGN luminosity is considered the primary driver of reverberation lag times. More luminous quasars tend to have larger time lags as observed by the relation (Bentz et al., 2013; Fonseca Alvarez et al., 2019), which matches basic photoionization expectations (). Furthermore, luminous quasars often exhibit lower amplitude variability (MacLeod et al., 2012; Vanden Berk et al., 2004), potentially reducing the likelihood of detecting a reliable lag measurement within short observing campaigns. To explore this connection, we adopt the luminosity as a measure of the quasar continuum luminosity. Whenever possible, we utilize directly measured values. For sources lacking direct measurements, we reconstruct them using , , and the bolometric corrections from Richards et al. (2006). Figure 1 illustrates the redshift and luminosity distribution of our targets. We performed a global comparison of the continuum luminosity within the gold samples of early SDSS-RM studies and the current 7-year lag measurements (Grier et al., 2017, 2019; Homayouni et al., 2020; Shen et al., 2024). This analysis did not reveal any statistically significant changes in the continuum luminosity distribution over the multi-year analysis. Additionally, we find no dependence between lag measurement success and source luminosity. The median continuum luminosity of the gold sample (in logarithmic scale) is , , for H, Mg II, and C IV, respectively, for the early SDSS-RM lag measurements, which is similar to the median range of , , for the parent sample of each targeted emission line. Here, the reported uncertainty is computed from the 16th and 84th percentiles. We find similar ranges for the gold-lag measurements in the 7-year results of Shen et al. (2024).
@@ 4.2 Rest-frame Equivalent Width
The rest-frame equivalent width (REW) of emission lines is another potential factor influencing successful lag measurement success. We investigate whether targets with larger REWs correlate with the most reliable lag measurements. For each emission line, we examine the potential relationship between its REW using PrepSpec outputs. REWs are measured from the continuum and flux provided by PrepSpec, utilizing either one-year, four-year, or seven-year average spectra. Figure 2 shows the comparisons between early-year and 7-year SDSS-RM data for the same emission lines, which reveal relatively consistent REWs over time. Furthermore, we find no direct correlation between the fraction of quasars with gold-lag measurements. Figure 2 depicts how the gold samples in each study are uniformly distributed within their respective parent samples, and the gold-lag success fraction does not reveal any specific trends with REW. While the one-year and four-year PrepSpec spectral fits did not incorporate any Fe II emission lines, the seven-year data included an Fe II template. Consequently, the 7-year measurements of Shen et al. 2024 (blue data points in Figure 2) appear to be systematically offset above the 1:1 line, suggesting that for all the three emission lines, the gold-lag estimates may be biased towards larger REWs. Comparison of the median REW values between the gold and parent samples in Table 2 indicates a higher success rate for the gold-lag sample in C IV and marginal for Mg II.
@@ 5.1 Fractional Variability
The results of this investigation are presented in Figure 4. We further compared the fractional rms variability between the early-year and 7-year studies for each emission line in the bottom panels of Figure 4. It is noteworthy that the 4-year and 7-year results largely follow a 1:1 trend. The deviation observed in the comparison between the 1-year and 7-year H fractional rms variability (bottom left corner of Figure 4) can be attributed to the variability of quasars on the observed-frame timescales. The damping timescale of quasars are longer than the seasonal monitoring duration.
@@ 5.3 Durbin-Watson Statistic
While our findings suggest the importance of variability as measured by SNR2, instances exist where a gold lag remains undetected despite relatively high SNR2 values. For time-series analysis software to effectively detect an RM lag, the light curve often requires a distinct “hook” or inflection feature in the continuum and line light curves. In this section, we investigate whether the presence of serial correlations in the continuum and emission line light curve would lead to a higher success in detecting a gold-lag measurement. Serial correlation refers to the dependence of serial data points, indicating a patterned behavior in the time series. A commonly used test for first-order autocorrelation, which assumes independent error terms, is the Durbin-Watson test (Durbin & Watson 1950). The Durbin–Watson test statistic is given by
@@
5.4 Logistic Regression
To further investigate which of the target properties most significantly influence the distribution of gold-lag measurements, we conducted a logistic regression for each of the six distinct sets of lag measurements. The independent variables considered in our model were , REW, fractional variability, SNR2 variability, and the continuum and line Durbin-Watson statistics. Our analysis of the early SDSS-RM lag measurements reveals that for the H lag measurements reported by Grier et al. (2017), the , along with both the continuum and line Durbin-Watson statistics, were the strongest predictors of a gold-lag measurement. However, a similar analysis for the Mg II lags in Homayouni et al. (2020) revealed only a marginal predictive power for (=0.048). For the CIV lags reported in Grier et al. (2019), the line Durbin-Watson statistic (=0.003) and, with less significance, the continuum Durbin-Watson statistic (=0.045) contributed to the identification of gold lags. For the 7-year lag measurements reported by Shen et al. (2024), our analysis indicates that for the H, as well as the C IV results, and the line Durbin-Watson statistic were identified as the strongest predictors. On the other hand, for the Mg II lags, the line Durbin-Watson statistic was identified as the single robust predictor of gold-lag measurements.
We also note that while the logistic regression reveals a negative correlation (as also illustrated in Figure 1), this trend may stem from inherent biases within our sample and the observational constraints of these studies. Specifically, the inverse connection to may reflect longer time delays in brighter sources (e.g., Bentz et al. 2013) that demand longer observational baselines not fully covered by these studies. With the exception of the Mg II line, the line Durbin-Watson statistic consistently emerged as the strongest predictor for identifying gold lags across all other emission lines examined in this work.
@@ 6.1 Cadence Intensity in SDSS-RM
Finally, we run PyROA for the simulated line light curves with their original continuum light curves and apply lag identification and alias removal process to get the final lag measurement. The reason we keep the original continuum light curves and do not remove epochs from them is that the photometric observations for continuum light curves are not as expensive to make as the spectroscopic observations and are made by telescopes worldwide, so the continuum light curves are therefore densely sampled. Specifically, we have 800 epochs along 11 years and 158 epochs in 2014 (the first year of line light curves) on average, which is extremely dense with a cadence of 3 days per epoch approximately. We also perform some tests on running PyROA for cadence reduced light curves for both continuum and emission lines. The distribution of interval between one epoch to the next is almost uniform for each year, i.e., we also reduced 20% of the total epochs for continuum light curves of several targets selected randomly from the gold sample and found similar lag PDFs using PYROA. Therefore, for convenience, removing epochs in the first year in the following text refers to randomly reducing the density of epochs in the first year of the original emission line light curve to 40% unless otherwise specified.
@@ 6.2 Impact of Reduced Cadence on Lag Success
We run all the simulated light curves through the same lag identification pipeline as Shen et al. (2024) and perform the following statistical analyses. We use two criteria to assess the difference between the original lag measurements and the simulated lag measurements for targets from the gold sample, and display the variation of simulated lags in Figure 8. The two criteria are:
@@
The lag selection criteria, specifically the , and and their change, are shown in the top panels of Figure 8. Most of the lag posterior samples given by PyROA will still have a single, prominent peak in the lower-cadence simulations, so we do not show the changes for the lag selection criteria . The top panel of Figure 8 displays the minimum statistical value across 50 simulations to capture the full range of lag selection criteria changes and resulting lag possibilities, potentially providing a better indication of the overall trend in a worst-case scenario. While the statistical values for and generally decreased in the lower-cadence simulations, most simulated lags still met the significance criteria. Figure 8 highlights several targets with original and values near the selection threshold that resulted in insignificant lag measurements under reduced cadence conditions. The simulated lags that were initially considered significant but later became insignificant are primarily found in cases with lower overall light curve quality. This suggests that the reduction in measurement quality during the first season pushed these lags below the original selection criteria. Despite having the lowest median observed-frame lag (31 days), the H sample still has the highest ratio of significant results. This may be attributed to the generally higher lag and values in the H sample, indicating its robustness to epoch reduction.
@@
We display the consistency level for significant results in the bottom row of Figure 8. As described above, the can quantitatively describe how well the original lag and the simulated lag are consistent within a given margin of error. We divided the sample into long- and short-lag groups based on their median original lags: 31 days for H, 155 days for Mg II, and 140 days for C IV. Short-lag targets exhibited a higher degree of consistency () compared to long-lag targets across all emission lines. The most pronounced difference was seen for H, with short-lag targets achieving a consistency rate, surpassing the of long-lag targets. Mg II and C IV also exhibited similar trends, with short-lag targets outperforming long-lag targets by small margins. Short Mg II lags showed a consistency rate of , compared to for long Mg II lags. Similarly, short C IV lags had a rate of , while long C IV lags were at . Among the three emission lines, C IV (with the longest median original lag) demonstrated the highest overall ratio of simulations that maintained consistent results within 1 errors, while H and Mg II exhibited lower ratios. The short-lag sample may exhibit higher consistency due to its smaller absolute changes in lags and associated errors compared to the long-lag sample. Overall, if we require the simulated lags are still statistically significant and consistent with original lags in , we can still recover 81%, 76%, and 86% of all simulations for H, Mg II, and for C IV respectively.
To inform future cadence design, we examined the relationship between lag recovery ratio, -band magnitude, and redshift, given that SDSS-RM targets are selected based on flux limits. Here, the lag recovery ratio is defined as the fraction of simulations in which the new lags obtained from simulations remain significant and consistent with the original lags within 1 (i.e., ) out of the total number of simulations. Figure 9 presents an overview of the lag recovery ratio binned by redshift (bin size = 0.4) and -band magnitude (bin size = 1). Overall, we observed a trend of higher success ratios for brighter -band magnitudes and lower redshifts. Several bins, particularly those near the boundary of the redshift-magnitude parameter space, contain fewer than five targets and may not provide statistically meaningful results. We therefore excluded these bins from Figure 9. Given our results based on seven years of observations from SDSS-RM, the limited number of significant lags detected at these redshifts and magnitudes, as also indicated by mock data in Shen et al. (2015a), underscores the challenges of measuring significant lags at these cosmological distances and magnitudes.
@@ 7 Summary and Implications for Future Industrial-Scale Surveys
The Durbin-Watson () statistic is a reliable indicator of gold-lag measurements, particularly when applied to emission line light curves. While the statistic for the continuum light curve shows a weaker correlation, analyzing the emission line reveals a stronger trend. For Mg II, the statistic exhibits a weaker correlation with gold lag success compared to H and C IV. This is likely due to Mg II’s lower variability amplitude (Sun et al., 2015). Overall, our analysis suggests that the statistic, especially when applied to emission line light curves, is a valuable tool for efficiently filtering targets and predicting the likelihood of obtaining reliable lag measurements, particularly in large datasets. Logistic regression further validates the dw emission line as a strong predictor but also indicates a link with continuum luminosity (see Section5.4). We caution, however, that the inverse correlation to source luminosity may be attributed to the longer time delays inherent in brighter sources (e.g., Bentz et al. 2013), potentially exceeding the observational baselines of current studies.
Additionally, we investigate the impact of reducing the SDSS-RM cadence on lag measurements based on the observed quasar light curves of (Shen et al., 2024). We simulate multi-year light curves with 40% reduction in cadence during the first year and compare with the existing lag measurements for the same target. While most statistically significant lags are preserved after cadence removal (90% or higher for all emission lines studied: H, Mg II, and C IV), a small percentage become insignificant. The consistency between original and simulated lags (within a margin of error) is generally higher for shorter lags compared to longer lags. Furthermore, we find that the recovery rate is slightly lower for fainter or higher redshift quasars (Figure 9). These results are consistent with earlier studies using shorter mock light curves Shen et al. (2015a). Overall, our simulations indicate that a mixed cadence strategy, featuring a period of higher cadence followed by lower cadence as in SDSS-RM, provides only a marginal improvement in gold-lag measurement compared to a more uniform cadence, which can still recover 90% of the lags.